# Low evidence for ringdown overtone in GW150914 when marginalizing over time and sky location uncertainty

### v3

@@ III Results and Discussion
Figure 2: Marginal posteriors on $A_{221}/A_{220}$ and the coalescence time as measured in the Hanford detector $t_{c}^{H}$; similar results are obtained for the Livingston detector (not shown). The horizontal axis shows the ratio between the $(2,2,1)$ and $(2,2,0)$ mode amplitudes. The vertical axis shows the difference between the measured $t^{H}_{c}$ and the reference time $t^{H}_{\rm ref}$, which is the maximum likelihood $t^{H}_{c}$ from Ref. [29]. The filled marginal histograms represent the 1D distributions for each parameter. The prior on $A_{221}/A_{220}$ is overlaid on the corresponding marginal posterior with a dotted black line. The 4-OGC $t^{H}_{c}$ posterior is overlaid on the $t_{c}$ marginal posterior in gray. Each point in the 2D scatterplot is colored according to its log likelihood value relative to the maximum log likelihood. The maximum likelihood point is denoted with a black cross. The 2D credible regions are indicated by the contour lines; shown are the 50th (dashed contour) and 90th (solid contours) percentiles. The Bayes factor for the $(2,2,1)$ mode is estimated from the $A_{221}/A_{220}$ marginal posterior using the Savagey-Dickey ratio. This is the ratio between the posterior density (solid black line) and the prior density (dotted black line) at $A_{221}/A_{220}=0$.
@@
Fig. 3 compares the sky location posteriors of our analysis with a full IMR analysis from 4-OGC [29]. Our analysis recovers most of the 90th percentile of the 4-OGC posterior. Our marginal posterior on coalescence time is also consistent with the full IMR analysis (see Fig. 2). This indicates that our sky and time marginalization implementation is able to recover nominal sky parameters for GW150914.
Figure 3: Sky location posteriors for our analysis depicted in Fig. 2 and a full IMR analysis from 4-OGC [29]. Contours indicate the 90th percentile of points for each run. The orange contour denotes the analysis conducted with sky and time marginalization described above. The blue contour denotes the 4-OGC analysis.
Figure 4: Final mass and spin posteriors for our analysis depicted in Fig. 2, a $(2,2,0)$-only ringdown analysis, and a full IMR analysis from 4-OGC [29]. Contours indicate the 90th percentile of points for each run. The blue contour denotes our ringdown analysis with both the $(2,2,0)$ and $(2,2,1)$ modes. The red contour denotes an analysis with only the dominant mode present in the ringdown. The black contour denotes the 4-OGC analysis. The maximum likelihood points of the models are indicated by the crosses. The median and 90% credible intervals from the overtone model are $71^{+30}_{-16}\,M_{\odot}$ and $0.70^{+0.25}_{-0.47}$ for the red-shifted mass and spin, respectively, while for the $(2,2,0)$-only model they are $84^{+13}_{-27}\,M_{\odot}$ and $0.84^{+0.09}_{-0.55}$.
We estimate the Bayes factor $\mathcal{B}$ between $Z_{220+221}$ and $Z_{220}$ by calculating the Savage-Dickey ratio [47] at $A_{221}/A_{220}=0$. Doing so, we obtain $\mathcal{B}=1.10\pm 0.03$, indicating low support in favor of the $(2,2,1)$ mode. Similarly, the median of the posterior distribution for $A_{221}$ deviates from 0 by $1.2\sigma$  [^3] The $1.2\sigma$ quoted here was computed using the relative amplitude posterior $A_{221}/A_{220}$. Recalculating using the absolute $A_{221}$ posterior yields the same result.. The reason for the small Bayes factor for the $(2,2,1)$ mode is evident from the marginal posterior plot in Fig. 2. While the maximum likelihood point is at $A_{221}/A_{220}=1.03$ and $t^{H}_{c}-t^{H}_{\rm ref}=0.03\,$ms, the $A_{221}/A_{220}$ posterior has support at zero over much of the $t^{H}_{c}$ posterior (similarly for $t^{L}_{c}$). It is only at earlier times that the posterior on $A_{221}$ is markedly shifted away from zero. This is consistent with previous studies which found the evidence for the $(2,2,1)$ mode increased at earlier times [3, 2, 26]. However, due to the relatively large timing uncertainty – our credible interval on $t^{H/L}_{c}$ spans roughly $5M$ – we cannot confidently claim detection of the overtone, as reflected by the Bayes factor. A signal with more SNR in the ringdown is needed.
@@
A detection of the $(3,3,0)$ sub-dominant QNM in the ringdown of GW190521 was claimed in Capano et al. [30] [^5] See also Siegel et al. [42], which found evidence for the $(2,1,0)$ and $(3,2,0)$ modes in the same event.. The sky location was also fixed in that analysis, and the Bayes factor was maximized over a grid of possible merger times. Our results here raise the question of whether the evidence for the sub-dominant mode in GW190521 would also decrease if the sky location and $t_{c}$ were marginalized over.
@@
Assuming the sky location used in Capano et al., the evidence for the $(3,3,0)$ mode stays elevated across a large range of times after merger [52]. In contrast, the evidence for the $(2,2,1)$ mode for GW150914 drops quickly after merger. A naive marginalization over the discrete times published in Capano et al. — accomplished by summing the log evidence over those times — yields a Bayes factor for the $(3,3,0)$ mode of 5000similar-toabsent5000$\sim$ 5000∼ 5000.666We report the Bayes factor for the $(3,3,0)$ mode as being $Z_{220{+}330}/\max(Z_{220},Z_{220{+}221})$, as was done in Capano et al. [30]. A reanalysis that uses a model for the pre- and post-merger signal and marginalizes over time and sky-location uncertainty is necessary to fully evaluate the evidence for the $(3,3,0)$ mode, and any other sub-dominant QNM, in GW190521. Nevertheless, our cursory analysis indicates that the evidence for the $(3,3,0)$ mode in GW190521 may in fact increase with full sky and time marginalization. We plan to revisit this in a future study.
@@
This research was conducted using PyCBC [57]. Data is available at [58].
Acknowledgements.
We wish to thank Gregorio Carullo and Juan Calderon Bustillo for providing helpful comments. A.C. was supported by funds from the Massachusetts Space Grant Consortium. C.C. acknowledges support from NSF award PHY-2309356. All computations were performed on Unity, a collaborative, multi-institutional high-performance computing cluster managed by UMass Amherst Research Computing and Data. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.

### v4

@@ III Results and Discussion
Figure 2: Marginal posteriors on $A_{221}/A_{220}$ and the coalescence time as measured in the Hanford detector $t_{c}^{H}$; similar results are obtained for the Livingston detector (not shown). The horizontal axis shows the ratio between the $(2,2,1)$ and $(2,2,0)$ mode amplitudes. The vertical axis shows the difference between the measured $t^{H}_{c}$ and the reference time $t^{H}_{\rm ref}$, which is the maximum likelihood $t^{H}_{c}$ from Ref. [29]. The filled marginal histograms represent the one-dimensional distributions for each parameter. The prior on $A_{221}/A_{220}$ is overlaid on the corresponding marginal posterior with a dotted black line. The 4-OGC $t^{H}_{c}$ posterior is overlaid on the $t_{c}$ marginal posterior in gray. Each point in the two-dimensional scatter plot is colored according to its log likelihood value relative to the maximum log likelihood. The maximum likelihood point is denoted with a black cross. The two-dimensional credible regions are indicated by the contour lines; shown are the 50th (dashed contour) and 90th (solid contours) percentiles. The Bayes factor for the $(2,2,1)$ mode is estimated from the $A_{221}/A_{220}$ marginal posterior using the Savagey-Dickey ratio. This is the ratio between the posterior density (solid black line) and the prior density (dotted black line) at $A_{221}/A_{220}=0$.
@@
Figure 3 compares the sky location posteriors of our analysis with a full IMR analysis from the fourth Open Gravitational-wave Catalog (4-OGC) [29]. Our analysis recovers most of the 90th percentile of the 4-OGC posterior. Our marginal posterior on coalescence time is also consistent with the full IMR analysis (see Fig. 2). This indicates that our sky and time marginalization implementation is able to recover nominal sky parameters for GW150914.
Figure 3: Sky location posteriors for our analysis depicted in Fig. 2 and a full IMR analysis from 4-OGC [29]. Contours indicate the 90th percentile of points for each run. The red contour denotes the analysis conducted with sky and time marginalization described in Methods. The black contour denotes the 4-OGC analysis.
Figure 4: Final mass and spin posteriors for our analysis depicted in Fig. 2, a $(2,2,0)$-only ringdown analysis, and a full IMR analysis from 4-OGC [29]. Contours indicate the 90th percentile of points for each run. The red contour denotes our ringdown analysis with both the $(2,2,0)$ and $(2,2,1)$ modes. The blue contour denotes an analysis with only the dominant mode present in the ringdown. The black contour denotes the 4-OGC analysis. The maximum likelihood points of the models are indicated by the crosses. The median and 90% credible intervals from the overtone model are $71^{+30}_{-16}\,M_{\odot}$ and $0.70^{+0.25}_{-0.47}$ for the red-shifted mass and spin, respectively, while for the $(2,2,0)$-only model they are $84^{+13}_{-27}\,M_{\odot}$ and $0.84^{+0.09}_{-0.55}$.
We estimate the Bayes factor $\mathcal{B}$ between $Z_{220+221}$ and $Z_{220}$ by calculating the Savage-Dickey ratio [47] at $A_{221}/A_{220}=0$. Doing so, we obtain $\mathcal{B}=1.10\pm 0.03$, indicating low support in favor of the $(2,2,1)$ mode. Similarly, the median of the posterior distribution for $A_{221}$ deviates from 0 by $1.2\sigma$  [^3] The $1.2\sigma$ quoted here was computed using the relative amplitude posterior $A_{221}/A_{220}$. Recalculating using the absolute $A_{221}$ posterior yields the same result.. The reason for the small Bayes factor for the $(2,2,1)$ mode is evident from the marginal posterior plot in Fig. 2. While the maximum likelihood point is at $A_{221}/A_{220}=1.03$ and $t^{H}_{c}-t^{H}_{\rm ref}=0.03\,$ms, the $A_{221}/A_{220}$ posterior has support at zero over much of the $t^{H}_{c}$ posterior (similarly for $t^{L}_{c}$). It is only at earlier times that the posterior on $A_{221}$ is markedly shifted away from zero. This is consistent with previous studies which found the evidence for the $(2,2,1)$ mode increased at earlier times [3, 2, 26]. However, due to the relatively large timing uncertainty—our credible interval on $t^{H/L}_{c}$ spans roughly $5M$—we cannot confidently claim detection of the overtone, as reflected by the Bayes factor. A signal with more SNR in the ringdown is needed.
@@
A detection of the $(3,3,0)$ subdominant QNM in the ringdown of GW190521 was claimed by Capano et al. [30] [^5] See also the work by Siegel et al. [42], which found evidence for the $(2,1,0)$ and $(3,2,0)$ modes in the same event.. The sky location was also fixed in that analysis, and the Bayes factor was maximized over a grid of possible merger times. Our results here raise the question of whether the evidence for the subdominant mode in GW190521 would also decrease if the sky location and $t_{c}$ were marginalized over.
@@
Assuming the sky location used by Capano et al., the evidence for the $(3,3,0)$ mode stays elevated across a large range of times after merger [52]. In contrast, the evidence for the $(2,2,1)$ mode for GW150914 drops quickly after merger. A naive marginalization over the discrete times published by Capano et al.–accomplished by summing the log evidence over those times—yields a Bayes factor for the $(3,3,0)$ mode of 5000similar-toabsent5000$\sim$ 5000∼ 5000.666We report the Bayes factor for the $(3,3,0)$ mode as being $Z_{220{+}330}/\max(Z_{220},Z_{220{+}221})$, as was done by Capano et al. [30]. A reanalysis that uses a model for the pre and postmerger signal and marginalizes over time and sky-location uncertainty is necessary to fully evaluate the evidence for the $(3,3,0)$ mode, and any other subdominant QNM, in GW190521. Nevertheless, our cursory analysis indicates that the evidence for the $(3,3,0)$ mode in GW190521 may in fact increase with full sky and time marginalization. We plan to revisit this in a future study.
@@
All computations were performed on Unity, a collaborative, multi-institutional high-performance computing cluster managed by UMass Amherst Research Computing and Data.
This research was conducted using PyCBC [57]. Data are available in Ref. [58].
IV acknowledgments
We wish to thank Gregorio Carullo and Juan Calderon Bustillo for providing helpful comments. A.C. was supported by funds from the Massachusetts Space Grant Consortium. C.C. acknowledges support from NSF Grant No. PHY-2309356.
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (888gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, and Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; and Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.